# CT2Rep: Automated Radiology Report Generation for 3D Medical Imaging

### v1

@@ Abstract
3D Medical Imaging, Chest CT Volumes, Radiology Text Reports, Report Generation, Longitudinal Data, Transformers
@@ 1 Introduction
Recognizing this gap, our work introduces CT2Rep, the first approach to automated radiology report generation for 3D medical imaging, specifically targeting chest CT volumes. CT2Rep leverages a novel 3D auto-regressive causal vision feature extractor, optimized for processing volumetric data. We also incorporate relational memory to utilize information from previous report generations, employing memory-driven conditional layer normalization to integrate this data into our framework. CT2Rep’s uniqueness, the first of its kind in 3D medical imaging, means that no directly comparable methods exist. Nonetheless, to demonstrate the effectiveness of our framework, we reasonably designed a baseline using a state-of-the-art vision encoder used for 3D chest CT volume interpretation, CT-Net [6], for report generation. CT2Rep outperforms this well-designed baseline method, showcasing the efficacy of our novel approach.
@@ 2.1 The Proposed Method
$r_{\text{out}}=\Phi_{\text{CT2Rep}}(x)=\Phi^{\text{transformer}}_{\text{dec}}(\Phi^{\text{transformer}}_{\text{enc}}(\Phi^{\text{visual}}_{\text{enc}}(x))).\vspace{-0.1cm}$
@@ 2.2 Longitudinal Data Utilization
Our dataset comprises 25,701 non-contrast 3D chest CT volumes from 21,314 unique patients, which expands to 49,138 volumes after applying multiple reconstructions tailored to different window settings [28]. Each volume features a resolution of $512\times 512$ pixels in the axial plane, with slice counts ranging from 100 to 600. The radiology reports associated with each volume are segmented into four sections: clinical information, technique, findings, and impression; however, only the findings section is utilized for report generation training. The same radiology report is used for each reconstructed volume of a single CT. The dataset is divided into a training set of 20,000 patients and a validation set of 1,314 patients, ensuring no overlap. CT volumes were converted to Hounsfield Units (HU) using slope and intercept values from the metadata and clipped to $[-1000~{}\text{HU},+200~{}\text{HU}]$ to represent the practical diagnostic limits of the HU scale [5]. Each volume was subsequently resized to achieve uniform spacing of 0.75 mm on the x and y axes and 1.5 mm on the z-axis. The volumes were either center-cropped or padded to achieve a consistent resolution of $(240)\times 480\times 480$.
@@ 3 Experiments and Results
Table 1: Quantitative evaluation showcases the effectiveness of our CT2Rep in comparison to a well-designed baseline and highlights how our enhanced method, CT2RepLong, leverages longitudinal data to improve performance.
@@ 3.3 Implementation Details
CT2Rep and the baseline method (see Sec. 3.1) were trained on 49,138 3D CT volumes and their corresponding reports (Sec. 2.3). We used the Adam optimizer with $\beta_{1}$ and $\beta_{2}$ hyperparameters set to 0.9 and 0.99, respectively. The learning rate was established at 0.00005 for the visual extractor and 0.0001 for the other parameters. A StepLR scheduler with a gamma of 0.1, a batch size of 1, and a maximum token count of 300 for the scheduler were employed. CT2RepLong and the ablation methods (Sec. 3.2) were trained on 28,441 pairs (Sec. 2.3), utilizing the same hyperparameters as CT2Rep. The training duration for all models was one week on a single NVIDIA A100 GPU, achieving 20 epochs. Inference takes approximately 35 seconds for CTRep and 50 seconds for CTRepLong.
@@ 4 Discussion and Conclusion
Acknowledgements. We would like to thank the Helmut Horten Foundation.

### v2

@@ Abstract
3D Medical Imaging, Chest CT Volume, Radiology Report, CT-RATE Dataset, Report Generation, Longitudinal, Transformers
@@ 1 Introduction
Recognizing this gap, our work introduces CT2Rep, the first approach to automated radiology report generation for 3D medical imaging, specifically targeting chest CT volumes. CT2Rep leverages a novel 3D auto-regressive causal vision feature extractor, optimized for processing volumetric data. We also incorporate relational memory to utilize information from previous report generations, employing memory-driven conditional layer normalization to integrate this data into our framework. To train our framework, we utilize the CT-RATE dataset [9], which consists of 25,692 non-contrast chest CT volumes, expanded to 50,188 through various reconstructions, from 21,304 unique patients, along with corresponding radiology reports. CT2Rep’s uniqueness, the first of its kind in 3D medical imaging, means that no directly comparable methods exist. Nonetheless, to demonstrate the effectiveness of our framework, we reasonably designed a baseline using a state-of-the-art vision encoder used for 3D chest CT volume interpretation, CT-Net [6], for report generation. CT2Rep outperforms this well-designed baseline method, showcasing the efficacy of our novel approach.
@@ 2.1 The Proposed Method
$r_{\text{out}}=\Phi_{\text{CT2Rep}}(x)=\Phi^{\text{transformer}}_{\text{dec}}(\Phi^{\text{transformer}}_{\text{enc}}(\Phi^{\text{visual}}_{\text{enc}}(x))).$
@@ 2.2 Longitudinal Data Utilization
We utilize 3D chest CT volumes along with corresponding radiology reports from the publicly available CT-RATE dataset [9]. For the development of CT2Rep, we employ all volumes and reports from the initial release of CT-RATE. Our dataset comprises 25,701 non-contrast 3D chest CT volumes from 21,314 unique patients, which expands to 49,138 volumes after applying multiple reconstructions tailored to different window settings [29]. Each volume features a resolution of $512\times 512$ pixels in the axial plane, with slice counts ranging from 100 to 600. The radiology reports associated with each volume are segmented into four sections: clinical information, technique, findings, and impression; however, only the findings section is utilized for report generation training. The same radiology report is used for each reconstructed volume of a single CT volume. The dataset is divided into a training set of 20,000 patients and a validation set of 1,314 patients, ensuring no overlap. CT volumes were converted to Hounsfield Units (HU) using slope and intercept values from the metadata and clipped to $[-1000~{}\text{HU},+200~{}\text{HU}]$ to represent the practical diagnostic limits of the HU scale [5]. Each volume was subsequently resized to achieve uniform spacing of 0.75 mm on the x and y axes and 1.5 mm on the z-axis. The volumes were either center-cropped or padded to achieve a consistent resolution of $(240)\times 480\times 480$.
@@ 3 Experiments and Results
Table 1: Quantitative evaluation showcases the effectiveness of our CT2Rep model in comparison to a well-designed baseline and highlights how our enhanced method, CT2RepLong, leverages longitudinal data to improve performance.
@@ 3.3 Implementation Details
CT2Rep and the baseline method (see Sec. 3.1) were trained on 49,138 3D CT volumes and their corresponding reports (Sec. 2.3). We used the Adam optimizer with $\beta_{1}$ and $\beta_{2}$ hyperparameters set to 0.9 and 0.99, respectively. The learning rate was established at 0.00005 for the visual extractor and 0.0001 for the other parameters. A StepLR scheduler with a gamma of 0.1, a batch size of 1, and a maximum token count of 300 for the scheduler were employed. CT2RepLong and the ablation methods (Sec. 3.2) were trained on 28,441 pairs (Sec. 2.3), utilizing the same hyperparameters as CT2Rep. The training duration for all models was one week on a single NVIDIA A100 GPU, achieving 20 epochs. Inference takes approximately 35 seconds for CT2Rep and 50 seconds for CT2RepLong.
@@ 4 Discussion and Conclusion
{credits}
4.0.1 Acknowledgements
We extend our gratitude to the Helmut Horten Foundation for their invaluable support of our research. Additionally, we would like to express our sincere appreciation to Istanbul Medipol University for providing the CT-RATE dataset.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.